OpenStudy (pottersheep):

Why can't x^1/3 = -2?

OpenStudy (anonymous):

It can. (-8)^(1/3) = -2

OpenStudy (jamesj):

It can. If (-2)^3 = -8 and hence (-8)^(1/3) = -2.

OpenStudy (pottersheep):

:S okay thanks

OpenStudy (ybarrap):

x^(1/3) = (-2) x = (-2)^3 = -8 so (-8)^(1/3) = -2 is possible.